This Question: 3 pts 15 of 46 (0 complete) This Test: 100 pts possible oe Find the critical values, XR andXL. forc0.95 and n- 12. A. 3.816 and 21.920 B. 3.053 and 24.725 C. 4.575 and 26.757 O D. 2.603 and 19.675
Answer #1

\mathbf{Solution}

A. 3.816 and 21.920

\mathbf{given:}c=0.95,n=12

1-\alpha =c

1-0 = 0.95

\alpha =0.05

\frac{\alpha }{2}=0.025

\left ( 1-\frac{\alpha }{2} \right )=0.975

\chi ^{2}_{R}=\chi ^{2}_{\left ( 1-\frac{\alpha }{2} \right ),n-1}=\chi ^{2}_{0.975,11}={\color{Red} 3.816}

\chi ^{2}_{L}=\chi ^{2}_{\frac{\alpha }{2},n-1}=\chi ^{2}_{0.025,11}={\color{Red} 21.920}